# LLM4SCREENLIT: Recommendations on Assessing the Performance of Large Language Models for Screening Literature in Systematic Reviews

### v1

@@ 3.1 Review Results
Comparing the 27 papers we included in this review with DC+ reveals significant methodological gaps in evaluating Gen-AI tools for systematic review screening, with an underutilization of appropriate performance metrics (see Figure˜2) and limited approaches to cost-benefit considerations:
@@
    Limited MCC adoption: Only 3 papers (11% of the sample) employed MCC, including DC+ and the two papers by Syriani et al. [39, 40], despite MCC’s advantages for handling imbalanced datasets common in systematic review screening. However, Syriani et al. rescaled MCC from its standard [1,1][-1,1] range to $[0,1]$, which may complicate cross-study comparisons. Failure to use MCC is a missed opportunity for robust performance evaluation, particularly given the inherent class imbalance in screening tasks.
@@
    Insufficient confusion matrix reporting: Only 6 papers (22%) reported complete confusion matrices, though 4 additional papers provided sufficient information (in terms of total positives, total negatives, sensitivity, and specificity) to enable reconstruction. This deficiency hampers meta-analysis and prevents readers from computing alternative metrics or conducting independent cost-benefit assessments tailored to their specific screening contexts.
@@
    • The prevalence of sensitivity/recall (22 papers, 78%) reflects the field’s appropriate concern with minimizing FNs (failing to identify relevant studies would undermine SR validity). However, this focus often occurred without complementary metrics to assess the trade-offs with overall efficiency.
    • The widespread use of accuracy (16 papers, 59%) alongside high sensitivity reporting suggests many researchers may not fully appreciate accuracy’s limitations for imbalanced datasets.
    • The very modest adoption of balanced accuracy (only 4 papers, 15%) and extremely low MCC usage indicates insufficient awareness of metrics specifically designed for class imbalance scenarios or the need for chance-anchoring.
    • Five papers used Kappa variants (some employing multiple versions), which is appropriate when there is no well-defined baseline. However, it is inappropriate to use standard performance metrics and kappa variants on the same confusion matrix unless you are comparing two potentially fallible classifications (e.g., classifications made by two LLMs, or a classification made by a single human researcher and an LLM). In addition, the PABAK variant does not seem to be a useful metric in any circumstances.
@@ 3.2 Additional Good Practices
(P3) Reporting performance metric variability: Six papers reported performance metric confidence intervals, standard errors, or standard deviations. However, there was little consistency in the methods used to calculate them. Two papers used properties of the binomial distribution, one used bootstrapping, and the others did not specify any particular method. Whenever using samples to make decisions about future LLM use, it is important to specify the variability of performance metrics in order to assess both whether performance is better than chance and whether there are significant differences between the performance of different LLMs.
(P4) Reporting LLM prompt output consistency & (P5) Reporting null outcomes: Two papers considered the internal consistency of LLMs, and three other papers reported the rate of null outcomes. This is necessary because LLMs can sometimes fail to deliver any output, can provide nonsensical outcomes, or can provide different outcomes for the same prompt. The rate of such problems should be reported, and the researchers should know how to calculate performance metrics fairly when such problems occur.
(P6) Investigating the reason for LLM misclassifications: Three papers assessed the reason for LLM classifications and found systematic problems with their prompts. Thus, when validating LLM performance with a view to using an LLM for an SR, analysing classification errors found during LLM validation can provide a means to refine and improve prompts.
@@
(P8) Addressing the risk of data contamination/leakage: In the context of LLM testing, there is always a risk that publicly available data used to train an LLM could be part of the data used to test LLM performance [21]. This is referred to as data leakage or data contamination. Only Tran et al.[41] reported that their retrospective study used data published after the LLM they studied was released. Dennstadt et al. [5], who used one prospective study and 10 benchmark studies, mentioned the issue, but did not suggest any solution. However, in general, there was little appreciation of the issue. None of the other seven papers that used data from public benchmarks mentioned the issue, and two papers that used retrospective studies that were not obtained from public benchmarks suggested that their data sets could be used as benchmarks.
(P9) Fairly assessing the costs of misclassification FNs vs. FPs: Nine papers (30%) explicitly considered differential costs of FNs vs FPs, but with substantial variations in their approaches:
@@ 3.3 Addressing Performance Metrics Limitations
A simple, working example (for two LLMs, llama3-Athene:70b and llama3.1:8b, from Table˜3) of how to calculate WMCC under class imbalance, with asymmetric costs reflected by a weight of $w=10$, is presented below.
Raw counts for llama3-Athene:70b LLM from Table˜1: TP=47,FN=125,TN=4242,FP=82TP=47,\,FN=125,\,TN=4242,\,FP=82. Assuming $w=10$, we may calculate WMCC for this LLM:
@@
This example shows the impact of weighting MCC. WMCC allows us to distinguish between llama3-Athene:70b and lama3.1:8b, which both had the same MCC value (see Table˜1), indicating that lama3.1:8b is a better classifier because $WMCC_{llama3.1:8b}^{w=10}>WMCC_{llama3-Athene:70b}^{w=10}$ due to lama3.1:8b having fewer FNs than llama3-Athene:70b, although it has substantially more FPs.

### v2

@@ 3.1 Review Results
The 29 papers (28 other papers and DC+) reveal significant methodological gaps in evaluating Gen-AI tools for systematic review screening, with an underutilization of appropriate unbiased performance metrics such as MCC (see Figure˜2) and limited approaches to cost-benefit considerations:
@@
    Limited MCC adoption: Only 3 papers (10% of the sample) employed MCC, including DC+ and the two papers by Syriani et al. [5, 6], despite MCC’s advantages for handling imbalanced datasets common in systematic review screening. However, Syriani et al. rescaled MCC from its standard [1,1][-1,1] range to $[0,1]$, which may complicate cross-study comparisons. Failure to use MCC is a missed opportunity for robust performance evaluation, particularly given the inherent class imbalance in screening tasks.
@@
    Insufficient confusion matrix reporting: After searching not only the papers but also any reported supplementary material, we found only 7 papers reported complete confusion matrices, though 4 additional papers provided sufficient information (in terms of total positives, total negatives, sensitivity, and specificity) to enable reconstruction. This deficiency hampers meta-analysis by preventing the construction of MCC, which is well-suited for meta-analysis (see [44]) and prevents readers from computing alternative metrics or conducting independent cost-benefit assessments tailored to their specific screening contexts. It should also be noted that reporting the proportions or percentages in each confusion metric element without reporting the total number of items also limits analysis, because without the raw counts, although MCC and WMCC can be calculated, a full cost-benefit analysis is impossible. As a further example of its importance, we also note that two of the 24 secondary studies in the SESR-Eval dataset [2] contain only included papers (zero excluded), meaning TN=0 and FP=0 by construction. For these studies, Precision is trivially 1.0 whenever any paper is correctly included, which inflates F1 and biases aggregate performance statistics—further illustrating why reporting complete confusion matrices is essential, as it makes such degenerate cases immediately visible.
@@
    • 22 papers used standard Recall/Sensitivity, while the additional 3 papers reported some form of recall-related metric, including weighted variants and partial metrics such as RecAT%, RecInc/RecExc, or Weighted Recall. While this prevalence is encouraging and reflects the field’s appropriate concern with minimizing FNs, the use of recall often occurred without complementary chance-anchored metrics (such as MCC) to assess the trade-offs with overall efficiency, and none of the five papers reporting workload savings incorporated FN costs into their calculations.
    • The widespread use of accuracy (17 papers) alongside high sensitivity reporting suggests many researchers may not fully appreciate accuracy’s limitations for imbalanced datasets.
    • The very modest adoption of balanced accuracy (only 4 papers, 14%) and extremely low MCC usage indicates insufficient awareness of metrics specifically designed for class imbalance scenarios or the need for chance-anchoring.
    • Seven papers used Kappa variants (some employing multiple versions), which is appropriate when there is no well-defined baseline. However, it is inappropriate to use standard performance metrics and kappa variants on the same confusion matrix because (with the exception of MCC) standard confusion matrix metrics assume that a gold standard exists, while kappa assumes that no gold standard exists. Kappa is appropriate when comparing two potentially fallible classifications (e.g., classifications made by a single human and/or one or more LLMs). In addition, the PABAK variant does not seem to be a useful metric in any circumstances.
@@ 3.2 Additional Good Practices
(P3) Reporting performance metric variability: When conducting a new SR, current SR standards (both medical [45] and SE [22]) mandate that any GenAI tool is validated and the validation is reported. Validation would need to be based on a sample of the papers to be screened. Researchers can then compare the performance of different LLMs with that obtained from human researchers. Confidence limits attached to the performance metrics are necessary to identify which LLMs (singly, or in combination with other LLMs or a single human, or also any other more deterministic tools) are most likely both to meet agreed performance levels and to introduce work load reductions of sufficient magnitude to balance potential loss of evidence [^2] In the context of retrospective studies based on existing SRs, the classification data is available for all the abstracts i.e., the population of abstracts for a specific SR have all been classified, so the standard error of any performance metric obtained from testing LLMs on the full data set is zero and calculating confidence intervals invalid within that SR; CIs remain valid when the benchmark uses subsampling or pools across SRs (see Section 3.5 and R5).. Seven papers reported performance metric confidence intervals, standard errors, or standard deviations, and five of the papers were undertaking studies based on samples. However, there was little consistency in the methods used to calculate confidence intervals. Two papers used properties of the binomial distribution, one used bootstrapping, and the others (apart from the paper by [2], which reported standard errors obtained from logarithmic regression) did not specify any particular method.
(P4) Reporting LLM prompt output consistency (optional) & (P5) Reporting null outcomes: Two papers considered the internal consistency of LLMs, and three other papers reported the rate of null outcomes. While detailed consistency analysis (varying temperature, top-k, etc.) may warrant separate investigation, basic reporting of output consistency is valuable for reproducibility. Null outcomes affect performance assessment regardless of whether they originate from the LLM itself or from infrastructure issues (e.g., network problems, API downtimes); both their rate and suspected cause should be reported, as they directly affect confusion matrix computation.
(P6) Investigating the reason for LLM misclassifications (optional, but recommended for iterative studies): Three papers assessed the reason for LLM classifications and found systematic problems with their prompts. This practice is particularly valuable for iterative studies where findings can inform prompt refinement. We acknowledge that it represents a shift from pure evaluation toward LLM improvement, and is therefore optional.
@@
(P8) Addressing the risk of data contamination/leakage: In the context of LLM testing, there is always a risk that publicly available data used to train an LLM could be part of the data used to test LLM performance [24]. This is referred to as data leakage or data contamination. Only Tran et al.[42] reported that their retrospective study used data published after the LLM they studied was released. Dennstadt et al. [8], who used one prospective study and 10 benchmark studies, mentioned the issue and pointed out that “To objectively assess how well an LLM-based solution can evaluate scientific publications for new research questions, large cultivated and independent prospective data sets on many different topics would be needed, which will be very challenging to create.” However, in general, there was little appreciation of the issue. None of the other seven papers that used data from public benchmarks mentioned the issue, and three papers that used retrospective studies that were not obtained from public benchmarks suggested that their data sets could be used as benchmarks.
(P9) Explicitly assessing and documenting the differential costs of FNs vs. FPs: None of the papers reporting work-saving metrics incorporated FN into their metric. They reported only the savings due to not requiring TNs to be processed and costs due to the additional processing of FPs. Failing to consider the risks posed by FNs is unlikely to lead to balanced assessments of the performance of different GenAI tools. In contrast, ten papers explicitly considered the differential costs of FNs vs FPs, but with substantial variations in their approaches:
@@
  • Huotala et al. [2] suggest setting a target of 95% for recall, with precision of approximately 50%.
@@ 3.3 Addressing Performance Metrics Limitations
Selecting the weight $w$ requires consideration of the specific SR context. We recommend: (1) stakeholder consultation to determine domain-specific consequences of missed evidence versus wasted screening effort; (2) sensitivity analysis to assess how different plausible values of $w$ affect model rankings; and (3) a default of $w=10$ as a reasonable starting point for many SR contexts, reflecting that missing a relevant study typically has greater consequences than including an irrelevant one for further review. The chosen weight should be documented and justified in the study protocol. For deployment studies with access to domain stakeholders, approach (1) is the primary method. For benchmarking studies without domain-specific stakeholders, sensitivity analysis (2) is the preferred approach as it reveals whether model rankings are robust to cost-ratio assumptions; the default w=10w{=}10 (3) can additionally be reported as a cross-study reference point. Empirical support for w=10w{=}10 as a conservative default comes from the SE-specific sensitivity analyses in Section˜3.4: the 12 SESR-Eval crossovers (median $w\approx 2.7$, max $w\approx 6.4$)—and the illustrative Syriani RL4SE crossover ($w\approx 6$)—all occur well below w=10w{=}10.
A simple, working example (for two LLMs, llama3-Athene:70b and llama3.1:8b, from Table˜1) of how to calculate WMCC under class imbalance, with asymmetric costs reflected by a weight of $w=10$, is presented below.
Raw counts for llama3-Athene:70b LLM from Table˜1: TP=47,FN=125,TN=$42$,FP=82TP=47,\,FN=125,\,TN=4242,\,FP=82. Assuming $w=10$, we may calculate WMCC for this LLM:
@@
This example shows the impact of weighting MCC. WMCC allows us to distinguish between llama3-Athene:70b and llama3.1:8b, which both had the same MCC value (see Table˜1), indicating that llama3.1:8b is a better classifier because $WMCC_{llama3.1:8b}^{w=10}>WMCC_{llama3-Athene:70b}^{w=10}$ due to llama3.1:8b having fewer FNs than llama3-Athene:70b, although it has substantially more FPs.
3.4 SE-specific Worked Examples
The preceding analysis uses confusion matrices from the biomedical DC+ study. To ground the same analysis in software engineering, we reuse the only SE paper in our review that reports complete confusion-matrix counts directly in the paper: Felizardo et al. [4] (ESEM’24), who evaluated ChatGPT-4 on two SE SRs at two classification thresholds (Likert $\geq$5 as the primary threshold and Likert $\geq$4 as a “conservative” threshold that favours inclusion). Table˜3 reports the raw counts from [4] together with Accuracy, Recall, Lost Evidence, MCC, and WMCC (w=10w{=}10) that we computed from those counts.
Confusion Matrix
SLR Threshold TP FP FN TN N Acc Recall Lost Ev. MCC WMCC
SLR1 Likert $\geq$ 4 50 35 14 35 134 63.4% 0.781 21.9% 0.292 0.195
SLR1 Likert $\geq$ 5 48 17 16 53 134 75.3% 0.750 25.0% 0.507 0.330
SLR2 Likert $\geq$ 4 128 68 20 232 448 80.3% 0.865 13.5% 0.605 0.557
SLR2 Likert $\geq$ 5 113 27 35 273 448 86.1% 0.764 23.6% 0.683 0.529
Table 3: SE-specific worked example: reanalysis of ChatGPT-4 screening results from Felizardo et al. [4] (ESEM’24) on two SE SLRs (using Felizardo et al.’s original dataset names). Confusion-matrix counts are taken verbatim from [4]; MCC and WMCC (w=10w{=}10) are computed by us. WMCC weights each positive example (TP, FN) ten times more than each negative example (TN, FP).
Four observations emerge from this SE-specific reanalysis:
  1. 1.
    Accuracy and MCC can prefer the wrong operating point. For SLR2, Accuracy favours the Likert $\geq$5 threshold (86.1% vs. 80.3%), and MCC agrees (0.683 vs. 0.605). Yet Recall (0.865 vs. 0.764), Lost Evidence (13.5% vs. 23.6%), and WMCC (0.557 vs. 0.529) all favour the more inclusive Likert $\geq$4 threshold. A practitioner who optimised on Accuracy or even on MCC would select the threshold that loses almost twice as much evidence (23.6% vs. 13.5%)—MCC corrects for chance but, because it treats FNs and FPs symmetrically, it still prefers the more specific threshold, whereas WMCC’s 10:1 FN weight correctly rewards the higher-Recall operating point.
  2. 2.
    WMCC can tolerate a small increase in Lost Evidence when workload reduction is large. For SLR1, the Likert $\geq$5 threshold has two more FNs (16 vs. 14) but eighteen fewer FPs (17 vs. 35) and eighteen more TNs (53 vs. 35) than Likert $\geq$4. Under a simple 10:1 weighted-cost model (FP + $10\cdotFN$), Likert $\geq$5 is marginally more costly (177 vs. 175), yet WMCC strongly prefers Likert $\geq$5 (0.330 vs. 0.195) because, unlike the simple weighted cost, chance-anchored metrics credit the large increase in correctly excluded abstracts (TN). WMCC therefore rewards screening-workload reductions that are large enough to compensate for a small increase in Lost Evidence—the cost asymmetry operates through the confusion matrix as a whole, not through FN counts alone.
  3. 3.
    MCC is chance-anchored; Accuracy is not. All four configurations show Accuracy between 63% and 86%, yet MCC ranges from 0.29 to 0.68—the chance-anchored metric reveals that the classifier’s improvement over random guessing is more modest than Accuracy suggests.
  4. 4.
    WMCC penalises FNs more severely than MCC. In every row, $\text{WMCC}(w{=}10)<\text{MCC}$, because the 10:1 FN penalty pulls the cost-sensitive score below the symmetric one—demonstrating that WMCC is doing what it advertises.
A nuance is that whilst Felizardo et al. report that only 2 of SLR1’s 16 FNs and 4 of SLR2’s 35 FNs ultimately remained as lost evidence after the full-text screening stage [4], this is simply a property of the retrospective SR pipeline (later stages can recover some FNs), not of the LLM-tool itself, and motivates the discussion of extending recommendations to later SR stages in Section˜4.3.
We also note that none of the SE studies reanalysed in this section (Felizardo et al., Syriani et al., Huotala et al.) address data contamination risks (P8/R8) since the SE papers being screened may have been in the LLMs’ training data. This differs from Tran et al. [42] who intentionally selected SRs published after the LLM’s training cutoff. We address this concern empirically below via a cross-domain reanalysis of DC+ (see Contamination-risk sensitivity analysis).
Beyond this worked example, the remaining SE papers in our review further illustrate the challenges and good practices we recommend. Note we include forward pointers to our researcher and practitioner recommendations detailed in Section 4.2 to assist cross-referencing for the reader.
  • Huotala et al. [1] (EASE’24): found their zero-shot GPT-3.5 and GPT-4 missed 35–50% of included papers when reproducing a prior SE SR screening—a direct SE-specific Lost Evidence illustration. Only Recall and Precision were reported so no confusion matrix, MCC metrics or cost-sensitive analysis.
  • Huotala et al. [2] (ESEM’25, SESR-Eval): This was the largest SE-specific benchmark (9 LLMs $\times$ 24 secondary studies, 34,528 primary studies). Accuracy ranged 0.34–0.85 and F1 ranged 0.07–0.92 across secondary studies. Two of the 24 secondary studies, [53] and [64] from their reference numbering, contain only included papers (I/E Ratio == 100:0 in their Table VI and Huotala et al. themselves note that [53] “contains only included studies”). This creates a degenerate case since trivially Precision=\,{=}\,TP/(TP++0)=\,{=}\,1.00 for any LLM that includes at least one paper, and the (TN+FP)(TN{+}FP) factor in the MCC denominator (Equation˜15) is zero, making MCC mathematically undefined. Including these two studies in per-study averages inflates aggregate Precision by ++0.05 and F1 by ++0.03 relative to the remaining 22 studies (computed from their Table IX) which is exactly the kind of artefact that complete confusion-matrix reporting (R4) makes visible. No MCC or cost-sensitive metrics were reported. We reanalyse the SESR-Eval data below.
  • Thode et al. [3] (IST): Reports only Recall and Precision for three LLMs, three prompt templates, and three SE datasets. A two-LLM ensemble reached 98–99% recall at 27% precision, however, no confusion matrix, no MCC, no cost-sensitive analysis, and no non-LLM baselines were reported. Data contamination risk was acknowledged but not mitigated, hence this is an SE example where our recommendations could have strengthened the evaluation.
  • Syriani et al. [5, 6]: These are the only SE papers in our review reporting MCC and comparing ChatGPT against non-LLM baselines (LR, RF, CNB, SVC via 5-fold CV with grid search), illustrating the SE operationalisation of P2/R9 when training data is available. However, they rescaled MCC from [1,1][-1,1] to $[0,1]$ using $\text{MCC}_{[0,1]}=0.5+\text{MCC}_{[-1,1]}/2$ [6], which inflates apparent performance and complicates cross-study comparisons. We reanalyse their data in Table˜4 below.
Syriani et al. reanalysis. Syriani et al. [6] report Recall, Specificity, and the number of included/excluded papers for each of five SE datasets (see their Table 1). For three datasets (UpdateCollabMDE, MobileMDE, MPM4CPS) we verified the confusion matrices directly against Syriani et al.’s public replication package [^4] Available at https://doi.org/10.5281/zenodo.10257742. The replication package also reveals that ChatGPT initially produced API errors (labelled “unknown”) for 15, 5, and 4 articles in these three datasets respectively. These were all retried and successfully resolved into INCLUDE/EXCLUDE decisions so the confusion matrices in Table 4 reflect the resolved outputs. The initial error rates (1.7%, 1.7%, 2.0%) are an SE-specific instance of the null-output issue addressed by R6.; the counts match exactly. For the remaining two datasets (RL4SE, DSMLCompo) the replication package contains results from a different experimental run, so we approximated the confusion matrices via $TP=\text{round}(\text{Recall}\times P)$ and $TN=\text{round}(\text{Specificity}\times N)$. Because these reported metrics are rounded to three decimal places, the reconstructed TP and TN counts may each differ from the true values by $\pm 1$, propagating to at most $\pm 0.007$ in MCC and $\pm 0.012$ in WMCC—well within the gap that separates classifiers in every comparison below. For the non-LLM baselines, Syriani et al. trained classifiers with 5-fold cross-validation repeated 10 times, so their reported Recall and Specificity are averages across folds and repeats. The confusion matrices we present for baselines are therefore synthetic (derived from averaged metrics) rather than observed from a single evaluation pass; we retain them because they are the only available basis for computing standard MCC and WMCC, but readers should note this caveat. Table˜4 shows the results for ChatGPT’s best prompt per dataset alongside the best non-LLM baseline, both selected on rescaled MCC as in [6] Table 5.
Confusion Matrix
Dataset Classifier TP FP FN TN Lost Ev. MCC WMCC
RL4SE ChatGPT 77 310 17 685 18.1% 0.298 0.511
(94+/995-) LR 56 130 38 865 40.4% 0.347 0.485
DSMLCompo ChatGPT 114 572 36 1961 24.0% 0.281 0.522
(150+/2533-) CNB 112 790 38 1743 25.3% 0.211 0.421
UpdateCollab. ChatGPT 51 286 6 532 10.5% 0.276 0.542
(57+/818-) SVC 27 123 30 695 52.6% 0.212 0.353
MobileMDE ChatGPT 47 50 8 187 14.5% 0.534 0.624
(55+/237-) SVC 40 45 15 192 27.3% 0.463 0.497
MPM4CPS ChatGPT 79 31 28 67 26.2% 0.423 0.256
(107+/98-) RF 35 18 72 80 67.3% 0.164 0.086
Table 4: Reanalysis of Syriani et al. [6]: ChatGPT (API model GPT-3.5-turbo-0613; best prompt per dataset) vs. best non-LLM baseline on five SE datasets. ChatGPT confusion matrices for UpdateCollabMDE, MobileMDE, and MPM4CPS are taken directly from the replication package; those for RL4SE and DSMLCompo are approximated from reported Recall, Specificity, and known class counts (shown in parentheses as positives+/negatives-), with counts accurate to $\pm 1$. Baseline confusion matrices are synthetic, derived from cross-validation-averaged metrics. MCC is on the standard [1,1][-1,1] scale; WMCC uses w=10w{=}10. Compare with Syriani et al.’s rescaled MCC[0,1]: ChatGPT ranged 0.638–0.767, baselines 0.584–0.734.
Three observations emerge from this reanalysis:
  1. 1.
    Rescaling inflates apparent performance. Syriani et al.’s rescaled MCC[0,1] for ChatGPT ranged 0.638–0.767 across the five datasets, suggesting moderate-to-good classification. Standard MCC[-1,1] ranges only 0.276–0.534—performance modestly above chance. A naïve cross-study comparison of Syriani’s rescaled values with, e.g., Felizardo’s standard MCC (Table˜3) would be seriously misleading, which is why R1 recommends reporting MCC unrescaled in [1,1][-1,1].
  2. 2.
    MCC and WMCC can disagree on the winner. For RL4SE, standard MCC favours LR over ChatGPT (0.347 vs. 0.298), but WMCC favours ChatGPT (0.511 vs. 0.485) because ChatGPT’s higher Recall (0.821 vs. 0.599) is rewarded by the 10:1 FN weight. A practitioner who chose the classifier using only MCC would select the one that loses 40% of relevant studies instead of 18%—precisely the scenario that cost-sensitive evaluation (R2) is designed to prevent. To verify that this ranking flip is not an artefact of the specific w=10w{=}10 choice, we computed WMCC for both classifiers across $w\in\{1,2,\ldots,20\}$: the crossover occurs at $w\approx 6$, meaning ChatGPT is the preferred classifier for any integer $w\geq 7$. The flip is therefore robust for any cost ratio that values missed evidence at least seven times more than wasted screening effort—a threshold most SR contexts would meet. (Because the RL4SE confusion matrices are approximated, we additionally confirmed that the flip persists across all $\text{TP}\pm 1$ and $\text{TN}\pm 1$ combinations for both classifiers at w=10w{=}10, that is 81/81 scenarios; see our replication-package script for details.) As the LR baseline confusion matrix is synthetic (derived from CV-averaged metrics), the RL4SE flip is best read as an illustrative single-pair example; the SESR-Eval reanalysis below provides the primary empirical demonstration using real per-article decisions.
  3. 3.
    Non-LLM baselines contextualise LLM performance. ChatGPT outperforms the best baseline on both MCC and WMCC in four of five datasets, but LR wins on MCC for RL4SE and all baselines show substantial Lost Evidence (25–67%). This illustrates R9, that when training data is available, baselines provide an informative reference point, but neither ChatGPT nor the baselines achieve levels of Lost Evidence that would support confident deployment.
[…]
Confusion Matrix
LLM (selection criterion) TP FP FN TN Acc Lost Ev. MCC WMCC
Claude 3.7 Sonnet (Acc-best) 189 213 326 8967 94.4% 63.3% 0.387 0.466
GPT-4.1 mini (MCC-best) 289 331 226 8849 94.3% 43.9% 0.481 0.604
GPT-4o (WMCC-best) 485 1743 30 7437 81.7% 5.8% 0.401 0.724
gpt-4.1-nano (Recall-best, excl. include-all) 505 7598 10 1582 21.5% 1.9% 0.093 0.228
Ministral 8B (include-all) 515 9180 0 0 5.3% 0.0% NaN NaN
Table 5: SESR-Eval reanalysis: Accuracy, MCC, and WMCC (w=10w{=}10) each select a different “best” LLM on the same 9,695-article SE secondary study (515 included, 9,180 excluded). Bold values mark the metric on which each model ranks first. The gpt-4.1-nano row shows the model a pure Recall-maximiser would select. Ministral 8B is included for contrast: it achieves 100% Recall by classifying every article as “include,” producing undefined MCC. Confusion matrices are computed from the raw per-article decisions in the SESR-Eval replication package [2].
Three observations emerge from this large-scale reanalysis:
  1. 1.
    Accuracy misleads at scale. Accuracy and MCC disagree on the best LLM in 11 of 22 evaluable secondary studies (50%; two studies with only included papers are excluded because MCC is undefined as explained above). In the example of Table˜5, the Accuracy-best model (Claude 3.7 Sonnet, 94.4%) loses 63.3% of the relevant evidence; even the MCC-best model (GPT-4.1 mini) still loses 43.9%. Only WMCC selects the model (GPT-4o) that retains 94.2% of the evidence—at the cost of additional FPs that increase human screening workload but do not compromise SR validity. The most extreme Accuracy-misleads instance occurs in a different study (3,703 articles, 45 included): GPT-4.1 mini achieves 97.5% Accuracy but loses 80% of the relevant evidence (MCC=\,{=}\,0.154).
  2. 2.
[…]
  3. 3.
    Triple metric disagreement demonstrates why MCC alone is insufficient. Table˜5 is, to our knowledge, the first empirical SE example in which Accuracy, MCC, and WMCC each point to a different winner. Accuracy is misleading under class imbalance (as argued throughout this paper); MCC corrects for chance but ignores asymmetric costs and still selects a model that loses 43.9% of evidence; only WMCC integrates chance-correction with cost asymmetry, selecting the model that preserves 94.2% of the evidence.
  4. 4.
    WMCC is not equivalent to Recall-maximisation. The same study’s gpt-4.1-nano has higher Recall (0.981 vs. 0.942) but lower WMCC (0.228 vs. 0.724) because it produces 7,598 FPs vs. GPT-4o’s 1,743 (Accuracy 21.5% vs. 81.7%). A pure Recall-maximiser would burden reviewers with $\sim$5,900 additional FPs to recover 0.04 of Recall; WMCC’s chance-anchored denominator correctly downgrades classifiers that buy Recall with excessive FPs.
Why MCC alone is insufficient and how WMCC resolves it: SE evidence from three reanalyses The problem. MCC is chance-anchored and imbalance-robust, but because it treats false negatives and false positives symmetrically, it can select the wrong classifier or operating point in SR screening. This is not a theoretical concern: in the Felizardo reanalysis, MCC agrees with Accuracy in favouring the threshold that loses more evidence; in the Syriani RL4SE data, MCC selects the classifier that loses 40% of relevant studies over one that loses 18%; and across the SESR-Eval dataset (9 LLMs $\times$ 24 SE secondary studies, 34,528 primary studies), MCC and WMCC disagree on the best LLM in 12 of 22 evaluable studies (55%). The solution. WMCC resolves this by encoding the FN:FP cost asymmetry directly into the chance-anchored correlation framework. In every disagreement, the WMCC-best model retains substantially more evidence than the MCC-best model. Sensitivity analysis shows that all ranking flips occur at modest cost ratios (median crossover $w\approx 2.7$, all below w=7w{=}7), meaning even a moderate preference for retaining evidence over saving screening effort is sufficient to change model selection—and empirically supporting w=10w{=}10 as a conservative default.
Contamination-risk sensitivity analysis.
A concern is that the SESR-Eval patterns could be contamination-driven: the 24 SE secondary studies are publicly indexed and may be in the cloud LLMs’ training corpora. As an orthogonal check, we reran the analyses on the complete confusion matrices from DC+ [7] (18 LLMs $\times$ 3 biomedical SRs; SI Figs. S1–S3), which uses different domains, different model families (locally served open-weight Llama/Gemma/Mistral/Mixtral/Qwen/Reflection, plus GPT-3.5/GPT-4o), and includes Review II (Neuro) whose underlying SR [46] was published in 2021 — before nearly every tested LLM’s training cutoff, making it the highest-contamination-risk SR in DC+.
The Accuracy-best LLM differs from the MCC-best LLM in all 3 DC+ SRs and in 26.3% of within-SR LLM pairs (101/384). Pairwise MCC/WMCC ranking flips occur in 31/384 pairs (8.1%), concentrated in Review II (28/153, 18.3%). If leakage caused the flip, we would expect the opposite pattern — fewer flips where models had “seen the answer”. The flip requires asymmetric error profiles (low FN at the cost of high FP), which leakage tends to erase rather than produce. This is strong evidence the flip is a property of the metrics, not an artefact of contamination. The llama3.1:8b vs llama3-Athene:70b comparison in Table 1 (DC+ Review I) is itself one such flip. Full results: dc_plus_reanalysis.R in the replication package.